Based on the following for the first six months:

Direct Labor Hours

Manufacturing Overhead

July

45,000

$295,000

August

60,000

$320,000

September

57,000

$323,000

October

52,000

$247,250

November

34,000

$178,200

December

25,000

$162,500

Required:

  1. Use the high-low method to estimate the cost formula
  2. Estimate the total overhead cost at an activity level of 48,000 machine hours, using the separate estimates you obtained for its components.
Answer #1
Required 1 :
Direct labor hours Manufacturing overhead
High level of activity [ August ] 60000 320000
(-) Low level of activity [ December ] 25000 162500
Difference 35000 157500
Variable cost per direct labor hour = Difference in Manufacturing overhead / Difference in direct labor hours = 157500 / 35000 4.5
Now let us calculate the fixed manufacturing overhead from the data of high level of activity
Total manufacturing cost = ( Variable cost per direct labor hours * Direct labor hours ) + Fixed manufacturing overhead
320000 = ( 60000 * 4.5 ) + Fixed manufacturing overhead
320000 = 270000 + Fixed manufacturing overhead
Fixed manufacturing overhead = 320000 - 270000 50000
Cost formula :
Y = 50000 + 4.5X
Required 2 :
Total overhead cost = 50000 + 4.5X
Total overhead cost = 50000 + 4.5*48000 266000